# Microphysical Origins of Gravitational Potentials in Astrophysical Systems

### v1

@@ III On the Observability of the Gravitational Constant $G$
As we saw, a subtle but significant aspect of gravitational physics is the fashion in which the gravitational constant $G$ enters observable phenomena. In many gravitational systems — particularly those that are virialized, such as stars, planets, and galaxies — quantities accessible to observation (e.g., gravitational redshift, line broadening, lensing deflection angles, escape velocities, and virial velocities) are determined by ratios or dimensionless combinations of parameters, wherein $G$ cancels out.
Similarly, in cosmology, the Friedmann equation in a spatially flat Universe can be cast in terms of dimensionless density parameters
$H^{2}(z)=H_{0}^{2}\left[\Omega_{r}(1+z)^{4}+\Omega_{m}(1+z)^{3}+\Omega_{\Lambda}\right],$ (3)
where $H(z)$ is the Hubble function, i.e. expansion rate of the Universe, $\Omega_{i}=\rho_{i}/\rho_{c}$, $\rho_{i}$ is the energy density of the i’th species, and $\rho_{c}=3H_{0}^{2}/(8\pi G)$ is the critical density. Here, $G$ is present only in the definition of the critical density; the expansion history, as determined solely by the ratios $\Omega_{i}$, is independent of $G$. Likewise, the comoving distance $d_{c}=\int_{0}^{z}\frac{dz^{\prime}}{H(z^{\prime})}$ is independent of $G$. All this implies that the angular diameter- and luminosity-distance are G-independent as well.
Only by combining cosmological measurements with an independent, non-gravitational determination of an absolute energy density—such as the CMB temperature (and thus the photon energy density)—can we directly constrain $G$. Without a direct measurement of the CMB temperature or the local baryon density, observations on cosmological scales would be insensitive to the empirical value of $G$.
It is only in the case of non-virialized systems, or in the direct interaction between two effectively pointlike masses (e.g., the Earth–Sun system), that $G$ appears explicitly in observable quantities, as in Newton gravitational law
$F=\frac{GMm}{r^{2}}.$ (4)
In these cases, the force or acceleration is directly proportional to $G$ and can be measured unambiguously.
@@
This pattern suggests that the role of $G$ is, in a sense, emergent: it is hidden in systems governed by self-gravity and revealed only when comparing gravitational effects to non-gravitational standards or when considering pointlike interactions. In this view, $G$ serves as a conversion factor between mass/energy and spacetime curvature, but its absolute value is often inaccessible in the absence of a fixed external scale. This observation highlights a profound feature of gravitational physics, with possible implications for our understanding of gravity as a fundamental versus emergent interaction.
IV Summary
@@
Typical astrophysical velocities are almost always non-relativistic, and the gravitational wells $\Phi$ characterizing these astrophysical systems are correspondingly shallow, i.e., $|\Phi|=r_{s}/R\ll 1$, where $r_{s}$ and $R$ are their typical Schwarzschild radius and actual size, respectively. According to the generalized Second Law of thermodynamics, this implies that the entropy budget of our Universe is much smaller than the holographic bound, $S_{dS}\sim 10^{122}$.
@@
This can be neatly explained by the fact that astrophysical systems are either regulated by cooling processes (governed by the electromagnetic interaction) or are required not to allow certain electromagnetic processes to occur. In other words, these systems must remain below specific temperature thresholds, otherwise their virial velocities would exceed their escape velocities and the systems would rapidly disperse. Typical temperatures are set by atomic or molecular physics, with $T\sim m_{e}\alpha_{e}^{2}$, so that virialized objects are typically characterized by $|\Phi|\sim\alpha_{e}^{2}\mu^{-\beta}$, where $\mu=m_{p}/m_{e}$ and $\beta$ is positive.
@@
Had $\alpha_{e}$ and $\mu$ been of order unity, the entire Universe would be teeming with BH’s and NS’s, with typical velocities approaching the speed of light, shortly after the Big Bang—regardless of its initial entropy state. So why is the Universe not predominantly populated by BH’s and NS’s ? Simply because $\alpha_{e}\ll 1$ and $\mu\gg 1$.
@@
While the reasoning laid out in this work is straightforward, the somewhat surprising conclusion—that although the existence of astrophysical systems depends on $G$ and specifically on its empirical value, certain properties characterizing these systems (e.g., gravitational redshift, line broadening, escape velocities, and gravitational lensing) are independent of $G$—appears to be overlooked in the literature. Specifically, the fact that large-scale velocities are determined solely by microphysics—essentially by the electromagnetic interaction rather than by gravitation—is not sufficiently appreciated. The main objective of this work was to emphasize this rather surprising fact. Although order-of-magnitude estimates of $M$, $R$ and $T$ for various astrophysical systems from first principles are discussed in the literature, the facts that (a) $\Phi$ is small due to the weakness of the electromagnetic interaction and the small electron-to-proton mass ratio, and (b) $\Phi$ is independent of $G$, seem to be overlooked, or at least not sufficiently acknowledged. In particular, we stress that this $G$-independence of $\Phi$ applies specifically to entire, virialized, naturally-formed astrophysical systems—such as stars, planets, and galaxies—that have formed and equilibrated through self-gravity and the relevant microphysical processes. Arbitrary non-virialized subsystems, like a piece cut from a planet, do not exhibit this property.
@@ Acknowledgments
This research was supported by a grant from the Joan and Irwin Jacobs donor-advised fund at the JCF (San Diego, CA).

### v2

@@ IV Implications of Extensive Planck Mass
As discussed above, a subtle but significant aspect of gravitational physics is the fashion in which the gravitational constant $G$ enters observable phenomena. In many gravitational systems – particularly those that are virialized, such as stars, planets, and galaxies – upper bounds on quantities accessible to observation (e.g., gravitational redshift, line broadening, lensing deflection angles, escape velocities, and virial velocities) are determined by microphysics with no reference to $G$. It is only in the case of non-virialized systems, or subsystems of otherwise virialized (or degeneracy pressure-supported systems) systems, or in the direct interaction between two effectively pointlike masses (e.g., the Earth–Sun system), that $G$ appears explicitly in observable quantities, as in Newton gravitational law. In these cases, the force or acceleration is directly proportional to $G$ and can be measured unambiguously.
@@
This pattern suggests that the role of $G$ is, in a sense, emergent; it is hidden in configurations where gravitation is balanced by other non-gravitational interactions and revealed only when comparing/measuring gravitational effects to/by non-gravitational standards or when considering pointlike interactions. This conclusion is ingrained and manifested throughout the analysis in Section III. In this context, $G$ serves as a conversion factor between energy density that is set by microphysics and the macroscopic spacetime curvature, but its absolute value is often inaccessible in the absence of a fixed external non-gravitational scale. This observation highlights a profound feature of gravitational physics, with possible implications for our understanding of the fundamental vs. emergent nature of gravity.
It should be stressed that the absolute value of $G$ is primarily determined by laboratory experiments on Earth, such as Cavendish-type measurements, and even then it is the measurement of gravitational interaction by comparison to non-gravitational forces that allows the empirical determination of $G$. Solar system observations – including lunar laser ranging, planetary orbit tracking, and spacecraft Doppler data – measure the product $GM$ (where $M$ is the mass of a celestial body) and cannot independently determine $G$ without knowing $M$. Therefore, these methods mainly constrain temporal or spatial variations of $G$ relative to the Earth-based value, rather than empirically inferring its absolute magnitude.
The following example illustrates one way in which $G$ might emerge. As discussed below Equation 6 – and shown with specific cases in Table 1 – systems whose characteristic density $\rho$ is fixed by microphysics (and does not involve $G$) exhibit scalings $R\propto G^{-\frac{1}{2}}$ and $M\propto G^{-\frac{3}{2}}$. By contrast, systems (like galaxies) whose formation process depends on $G$ have $\rho\propto G$, leading instead to $R\propto G^{-1}$ and $M\propto G^{-2}$ as required by the $G$-independence of $\Phi$.
Here we focus on the more generic case where $\rho$ truly does not depend on $G$, and in particular on habitable planets. Our objective is to explain – at the dimensional analysis level – why terrestrial Cavendish-type experiments measure an effective Planck mass $m_{pl}\sim 10^{19}$ GeV rather than some other value. From Table 1, one finds that the typical radius of such a planet is $R\sim\frac{\sqrt{\epsilon\alpha_{e}}m_{pl}}{m_{p}}a_{0}$ where $a_{0}$ is the Bohr radius, $\alpha_{e}$ the fine-structure constant, and $\epsilon$ a parameter of order $O(0.1)$ as was already defined in Section II but we reiterate for convenience. Since atoms in a rocky planet are tightly packed, we also have $R\sim(2a_{0})N^{1/3}$ where $N$ is the number of atoms. Combining these two estimates to eliminate $R$ we arrive at $m_{pl}=\frac{2m_{p}}{\sqrt{\epsilon\alpha_{e}}}N^{1/3}$.
Conventionally, we would choose $k=1$ as defined at the beginning of this section, i.e. we set the gravitational mass equal to the inertial mass thereby fixing $G$, and interpret this relation as saying that the number of atoms $N$ in a habitable planet is determined by $m_{pl}$, $m_{p}$, $\alpha_{e}$ and $\epsilon$. Indeed, a similar approach has been adopted in the context of estimating the number of protons in a typical nuclear burning star, e.g. [30, 20, 31]. However, adopting the viewpoint that $m_{pl}$ is merely an emergent conversion factor, and in addition noting that empirically $\tfrac{2\,m_{p}}{\sqrt{\epsilon\,\alpha_{e}}}$ is of order the Higgs mass $m_{H}$, this logic can be reversed, and we are led to tentatively conjecture instead that
$m_{\rm pl}\sim m_{H}\,N^{1/3},$ (13)
i.e. that $m_{\rm pl}$ is determined by $N$ and not the other way around – the larger $N$ is the weaker is gravity, and the larger is $m_{pl}$. From this perspective, the Earth-based laboratory determination of a small $G$ (equivalently a large $m_{pl}$) simply reflects the fact that typical habitable planets like Earth contain $N\sim\epsilon^{\frac{3}{2}}\lambda^{\frac{3}{2}}\sim 10^{51}$ atoms, producing a 17$\sim$ 17∼ 17-order-of-magnitude hierarchy $m_{\rm Pl}/m_{H}\sim N^{1/3}$. Following this logic and repeating Cavendish-like experiments on, e.g. the moon or Mars (while fixing $k=1$) will result in different inferred values for $G$ (i.e. values different from that inferred on Earth) if our assumption that $G$ is an emergent value following Equation 13 rather than a universal constant is correct. This is a falsifiable prediction of our proposal. It should be also emphasized that the proposed Planck-to-Higgs mass relation, Equation 13, is not rigorously derived from any first princples and it may well be an empirical coincidence that it empirically fits Earth-based measurements. However, the scaling $m_{pl}\propto N^{1/3}$ is more robust as was just discussed above Equation 13, and the proprtionality factor can be some other non-Higgs scale. Nevertheless, it would be a remarkable coincidence if the latter just happens to nearly coincide with $m_{H}$.
Even the number of particles in Earth is determined by fundamental principles to be $N\sim\epsilon\lambda^{\frac{3}{2}}$. Fitting the order-of-magnitude Earth compactness $|\Phi|\sim\epsilon\alpha_{e}^{2}\mu^{-1}$ (as shown in Table 1) to its measured compactness we obtain that $\epsilon\sim 0.05$. Combining the proposed scaling $m_{pl}\propto N^{1/3}m_{H}$ with $N\sim\epsilon\lambda^{\frac{3}{2}}$ we than conclude that $m_{p}\sim\sqrt{\epsilon\alpha_{e}}m_{H}$. Assuming $\epsilon\sim 0.05$ and the measured proton mass we remarkably obtain the order-of-magnitude estimate $m_{H}\sim 50$ GeV. Crucially, within each system, the ratio $m/m_{pl}$ is independent of $N$ because both gravitational mass and Planck mass scale identically with $N^{1/3}$. This ensures gravitational interactions between any two systems with different particle numbers remain universal and consistent with the equivalence principle, as the gravitational coupling between two systems with masses $m_{1}$ and $m_{2}$ depends only on these ratios as $\alpha_{g}^{1,2}=Gm_{1}m_{2}=\frac{m_{1}}{m_{pl}}\frac{m_{2}}{m_{pl}}$.
Thinking in terms of current terrestrial experimental standards and precision of $G$-inference could be disheartening. Indeed, conducting a Cavendish-like experiment on, e.g. the moon or Mars is a daunting task that presents both intriguing scientific opportunities and significant challenges. Although technically demanding, successfully conducting this experiment could provide significant insights into the nature of gravity, making it a scientifically valuable endeavor worth pursuing despite the obstacles. This is especially true given that falsifying either the universality of $m_{pl}$ or its proposed $\propto N^{\frac{1}{3}}$ dependence on, e.g., the moon is not as demanding as a full-fledged precision inference of $G$ with current standards of precision measurements. Specifically, according to Equation 13 the fractional relative divergence of $m_{pl}$ inferred in two systems that contain $N_{1}$ and $N_{2}$ particles respectively is expected to be $\Delta m_{pl}/m_{pl}=1-(N_{2}/N_{1})^{\frac{1}{3}}$. With the moon-to-Earth mass ratio of 0.012$\sim$ 0.012∼ 0.012 and assuming small binding energies in both systems we obtain $\Delta m_{pl}/m_{pl}\sim 0.77$, and consequently a very modest 10$\sim$ 10∼ 10 percent precision is already more than sufficient to falsify the proposed scaling.
Even on Earth $G$ can vary from place to place on its surface, because the surface of planets is shaped not only by gravity but also by the electromagnetic interaction as well, i.e. by the strength of matter as measured by its bulk modulus (compressibility). On planets, this is typically of order $Ry/(2a_{0})^{3}$, e.g. [23]. The estimated mass of Mount Everest, approximated as a cone with a base radius of about 5 km and height 8.8 km, and assuming an average Earth density of $5gr/cm^{3}$, is $\sim 10^{18}gr$. Comparing this to Earth mass $5.97\times 10^{27}gr$ and assuming similar mass densities for both we find $\Delta N/N=O(10^{-10})$, the fractional Planck mass difference between sea level and the summit of mount Everest is then $\Delta m_{pl}/m_{pl}\approx\Delta N/(3N)$ which is smaller than one part in ten billion. This is much smaller a difference than current precision in $G$ measurements. The latter is not much better than one part in $10^{5}$. Thus, whereas a moon-based Cavendish-like experiment can put Equation 13 to test, near-future Earth-based falsification of this conjecture seem to be not feasible.
One immediate consequence of Equation 13 is that the Higgs-to-Planck energy-density ratio becomes
$\rho_{vac}\sim m_{pl}^{4}\propto N^{4/3}.$ (14)
On cosmological scales where $N=O(10^{88})$ is dominated by CMB photons and primordial neutrinos this ratio is $O(10^{120})$. This is tentatively close to the $\sim$122 orders-of-magnitude fine tuning of the observed $\rho_{\Lambda}$, i.e. the energy density associated with the cosmological constant, and the naive expectation for the zero-point vacuum energy density, $\rho_{vac}=O(m_{pl}^{4})$ [16]. From the present work perspective, and as we explicitly saw in Section III, the cosmological description of a cosmic evolution dominated by $w=-1$, i.e. a vacuum energy-like cosmological constant, is in the microscopic, intensive, frame that corresponds to $N=1$. Compared to the extensive frame description with $N=O(10^{88})$, the 122$\sim$ 122∼ 122 orders of magnitude gap between observations and naive theoretical expectations can be almost fully bridged by realizing that the apparent mismatch results from comparing intensive to extensive quantities in a very complex system that contains $O(10^{88})$ particles.
Although the extensive vs. intensive perspective requires a markedly unusual treatment of $G$, that admittedly diverges even from other emergent-gravity perspectives, it offers a new perspective – by virtue of its $N$-dependence – on both the Planck-Higgs mass hierarchy and the ‘unexpectedly’ small cosmological constant.
At first glance, it may seem that we have simply traded the Higgs-Planck mass hierarchy and the cosmological constant problems by the question; why is our Universe so vast in complexity, i.e. why do habitable planets, nuclear-burning stars, and the observable Universe contain $N\sim 10^{51}$, $10^{57}$ and $10^{88}$ particles, respectively? However, our interpretation offers a unified, quantitative scaling of $G$ with system complexity $N$, thereby bringing both hierarchies under one umbrella. The ultimate (qualitative, if not quantitative) resolution of “why so large complexity?” may ultimately – perhaps unsurprisingly – be anthropic; only a sufficiently complex cosmos can give rise to complex observers.
V Summary
@@
Typical astrophysical velocities are almost always non-relativistic, and the gravitational wells $\Phi$ characterizing these astrophysical systems are correspondingly shallow, i.e., $|\Phi|=r_{s}/R\ll 1$, where $r_{s}$ and $R$ are their typical Schwarzschild radius and actual size, respectively. According to the generalized Second Law of thermodynamics, this implies that the entropy budget of our Universe is much smaller than the holographic entropy bound, $S\sim 10^{122}$ [11].
@@
This can be neatly explained by the fact that astrophysical systems are either regulated by cooling processes (governed by the electromagnetic interaction) or are required not to allow certain electromagnetic processes to occur. In other words, these systems must remain below specific temperature thresholds, otherwise their virial velocities would exceed their escape velocities and the systems would rapidly disperse. Typical temperatures are set by atomic or molecular physics, with $T\sim m_{e}\alpha_{e}^{2}$, so that virialized objects are typically characterized by $|\Phi|\sim\alpha_{e}^{2}\mu^{-\gamma}$, where $\mu=m_{p}/m_{e}$ and $\gamma$ is positive.
@@
Had $\alpha_{e}$ and $\mu$ been of order unity, the entire Universe would be teeming with BH’s and NS’s, with typical velocities approaching the speed of light, shortly after the Big Bang – regardless of its initial entropy state. So why is the Universe not predominantly populated by BH’s and NS’s? Simply because $\alpha_{e}\ll 1$ and $\mu\gg 1$.
@@
Whereas the reasoning laid out in this work is straightforward, the somewhat surprising conclusion – that although the existence of astrophysical systems depends on $G$ and specifically on its empirical value, large-scale velocities at the system boundary are determined solely by microphysics – essentially by the electromagnetic interaction rather than by gravitation – is not sufficiently appreciated in the literature. One objective of this work was to emphasize this rather surprising fact. Although order-of-magnitude estimates of $M$, $R$ and $T$ for various astrophysical systems from first principles are discussed in the literature, the facts that $|\Phi|$ is small due to the weakness of the electromagnetic interaction and the small electron-to-proton mass ratio, and $|\Phi|$ is independent of $G$, seem to be overlooked, or at least not sufficiently acknowledged. In particular, we stress that this $G$-independence of $|\Phi|$ applies specifically to entire, virialized, naturally-formed astrophysical systems – such as stars, planets, and galaxies – that have formed and equilibrated through self-gravity and the relevant microphysical processes. Arbitrary non-virialized subsystems, like a piece cut from a planet, do not exhibit this property.
In gravitational systems, distinguishing between intensive and extensive quantities offers a coherent framework to tentatively address the mass hierarchy and dark energy problems. Specifically, the inertial mass scale $m_{H}$, set by particle physics, is a constant of nature, independent of system size or particle number $N$. In contrast, we put forward the proposal that the gravitational mass scale, characterized by the effective Planck mass $m_{pl}$, scales as $m_{pl}\sim m_{H}N^{1/3}$ (ignoring possible geometric factors – or others – of order unity), a conjecture that could be directly falsified by future moderate-precision moon-based Cavendish-like experiments. For example, Earth’s particle content $N\sim 10^{51}$ quantitatively explains the observed hierarchy $m_{pl}/m_{H}\sim 10^{17}$ measured in Cavendish experiments, linking microscopic particle masses to macroscopic gravitational scales. Extending this to cosmology, vacuum energy density $\rho_{vac}$ is measured as we saw in Section III as an intensive quantity, i.e. $\rho_{vac}\propto m_{pl}^{4}\propto N^{4/3}$. With $N\sim 10^{88}$, approximately the number of CMB photons in the observable Universe, this scaling naturally suppresses naive particle physics predictions by about 120 orders of magnitude; observing the Universe from within, cosmological measurements probe intensive quantities that reflect the gravitational scaling tied to the Universe particle content, e.g. the energy density rather than total mass of the Universe (much like the energy density of the CMB does not depend on the total number of CMB photons in the Universe and the mass densities of e.g., planets, WD and NS depend exclusively on microphysics). As a consequence, the ‘naturalness’ problems associated with the mass hierarchy and cosmological constant could be rephrased as why is our Universe so complex, i.e. why, e.g. planet Earth and the observable Universe contain $O(10^{51})$ and $O(10^{88})$ particles, respectively. This question is rarely asked in the context of these puzzling problems.
@@ Acknowledgments
Anonymous referees are acknowledged for their constructive comments that stimulated a significant revision and expansion of the scope of this work. This research was supported by a grant from the Joan and Irwin Jacobs donor-advised fund at the JCF (San Diego, CA).